OpenStudy (anonymous):

With the height 8 m and the volume 150(pi) m^3, what is the cone's diameter? a) 7.5 m b) 5√3 m c) 7.5√3 m d) 15 m @Whiterabbit541 whatcha think? I think it's B) 5√3 m

OpenStudy (anonymous):

volume of a cone = (1/3)(pi) (r^2) (h) 150(pi) = (1/3)(pi) (r^2) (8) 150(pi)/(pi) = (1/3)(pi)/(pi) (r^2) (8) 150 = (1/3) (r^2) (8) 150 x 3 = (1/3 x 3/1) (r^2) (8) 450 = (r^2) (8) 450/ 8 = (r^2) (8)/8 56.25 = (r^2) 7.5 = r 7.5 x 2 = 15 diameter = 15 m answer = d) 15 m

OpenStudy (anonymous):

@Whiterabbit541 are you completely sure?

OpenStudy (anonymous):

yes :3